The KLM company started its operations in 2015. Three years later, the balance sheet for December 31, 2017. showed the following account balances (there were no other accounts listed):

Cash                                         ?

Accounts receivables                   $129,000

Loan made to the ABC company    $35,000

Accumulated depreciation             $50,000

common stock                            $500,000

Accounts payable                        $161,000

Wages payable                           $4,000

PP&E                                        $200,000

Retained earnings                        $304,000

Inventory                                   $49,000

Prepaid rent                                $12,000

During 2018 the following transactions occurred:

1. Collections from customers were $753,000

2. A machine (PP&E) was purchased for $250,000 all in cash.

3. The total cost of goods sold during the year was $656,000

4. Sales, all on credit, were $777,000

5. Marketing expenses were $25,000 of this amount, $18,000 was paid in cash.

6. The CEO of KLM announced she would resign her position effective June 30, 2018. Market price dropped $1.50 a share.

7. KLM purchased $666,000 worth of inventory, on account.

8. The employees of KLM earned $48,000 as wages. Cash payments to employees were $42,000

9. Payments on accounts payable were $777,000

10. Depreciation expense was $30,000

11. Rent in the amount of $30,000 was paid, in cash, during the year. The “Prepaid rent” account balance at the end of the year was $3,000

12. The ABC company was supposed to pay KLM $5 as interest for 2017 on its loan; however, due to a strike in the postal service the payment had arrived on January 15, 2018.

13. A dividend of $48,000 was declared during 2017, but was not paid.

14. A long-term loan in the amount of $50 was taken from a local bank.

Required

a. What was the balance of the Cash account on December 31, 2018?

b. Show the effect of each transaction during 2018 on the company’s accounts

c. Compute the company’s net income for fiscal 2018

d. Compute the balance in the company’s retained earnings account as of December 31, 2018.

e. Make a ending balance sheet for fiscal year 2018.
Answer #1
Trial Balance:
Req b) Effect of eash transaction during 2018 in accounting equation:
Assets = Liabilities Stockholders' Equity
Cash AR Inven. Prep. Rent Loan to ABC PPE Acc Dep = AP WP Divi Pay. LT Loan CS RE
Op. Bal. 594000 129000 49000 12000 35000 200000 -50000 = 161000 4000 500000 304000
1 753000 -753000 =
2 -250000 250000 =
3 -656000 = -656000
4 777000 = 777000
5 -18000 = 7000 -25000
6-No entry =
7 666000 = 666000
8 -42000 = 6000 -48000
9 -777000 = -777000
10 -30000 = -30000
11 -30000 -9000 = -39000
12 5 = 5
13 = 48000 -48000
14 50 = 50
Balance 230055 153000 59000 3000 35000 450000 -80000 = 57000 10000 48000 50 500000 235005
850055 = 850055
Req a) Cash account balance on Dec 31, 2018 is $230055.
Req c) Amount $
Sales rev. 777000
Int. rev. 5
total Revenue 777005
Less:COGS 656000
Gross Profit 121005
Less:Expenses:
Marketing -25000
wages -48000
Depreciation -30000
Rent -39000
total expenses -142000
Net Loss -20995
The company's net income for fiscal 2018 = -20995
Req d) Retained earnings as on Dec 31, 2018 is $235005, detailed:
Opening balance 304000
Less:Net Loss -20995
Less:Dividend declared -48000
Closing balance 235005
Req e)
Balance Sheet:
Assets; Amount $
Cash 230055
AR 153000
Inventory 59000
Prep. Rent 3000
Current assets 445055
Investment 35000
PPE 450000
Acc Dep. -80000
Total Assets 850055
Liabilities Amount $
AP 57000
Wage Pay 10000
Divi. Pay 48000
Current Liabilities 115000
Com Stock 500000
Ret Earn. 235005
Total Liabilities 850005